Calculate the Delta G for the following scenarios
Answer #1

Delta G = RT ln [X]in/[X]out

= 8.314 x 310 ln 30/1

Delta G = +8766 J/mol

Delta G = RT ln [X]in/[X]out

= 8.314 x 310 ln 30

= +8766 J/mol

delta due to membrane potential

ΔG = (z)(F)(Vm) where

z= the charge on the ion (+1 in this case)

F= 96485

Vm= the membrane potential, about − 60mV

delta G = +1 x 96485 x 0.06

deltaG = -5789 J

So total delta G = 8766 J - 5789 J = +2977 J/mol

delta G = RT ln [X]in/[X]out

= 8.314 x 298 ln 12 x10-3/145 x 10-3

=-6173 J/mol

ΔG = (z)(F)(Vm)

delta G = -1 x 96485 x -0.09

delta G = +8683 J

total delta G = 8683 - 6173 = +2509 J/mol

at pH 8; [H+] = 10-pH = 10-8 = 1 x 10-8 M

at pH 7; [H+] = 10-pH = 10-7 = 1 x 10-7 M

delta G = RT ln [X]in/[X]out

= 8.314 x 298 ln 1 x10-8/1 x 10-7

=-5704 J/mol

ΔG = (z)(F)(Vm)

delta G = +1 x 96485 x -0.18

delta G = -17367 J

total delta G = -17367 - 5704 = -23072 J/mol

SO only the last process is spontaneous and all the other require energy to happen